### v2

@@ 3.4 Training and Inference
$\mathcal{L}=\lambda D+R=\lambda\sum_{v=1}^{2}d(\boldsymbol{x}_{v},\boldsymbol{\hat{x}}_{v})+\sum_{v=1}^{2}\big{(}R(\boldsymbol{\hat{y}}_{v})+R(\boldsymbol{\hat{z}}_{v})\big{)},$ (7)
where $d(\boldsymbol{x}_{v},\boldsymbol{\hat{x}}_{v})$ represents the distortion between the original image $\boldsymbol{x}_{v}$ and reconstructed image $\boldsymbol{\hat{x}}_{v}$ under a given metric such as MSE or MS-SSIM [43]. $R(\boldsymbol{\hat{y}}_{v})$ denotes the estimated bitrate of the quantized representation $\boldsymbol{\hat{y}}_{v}$ and $R(\boldsymbol{\hat{z}}_{v})$ denotes the estimated bitrate of the hyper latent $\boldsymbol{\hat{z}}_{v}$. $\lambda$ is a hyper parameter that is used to balance the compression rate $R$ and distortion $D$.
4 Experiments
4.1 Experimental Setup
Dataset. We evaluate the coding efficiency of our proposed CAMSIC on two public stereo image datasets:
  1. (1)
    Cityscapes [17]: A dataset of urban outdoor scenes with distant views. It includes 5000 image pairs at a resolution of 2048$\times$1024 pixels. We divide these into 2975 pairs for training, 500 pairs for validation, and the remaining 1525 pairs for testing.
  2. (2)
    InStereo2K [7]: A dataset of indoor scenes with close views. It contains 2060 image pairs at a resolution of 1080$\times$860 pixels. We allocate 2010 and 50 stereo image pairs for training and testing, respectively.
We randomly crop the images into $256\times 256$ patches for data augmentation during training, and evaluate different codecs on full-resolution images. When the resolution of the source image is not supported, we apply the replicate padding to the right and bottom sides of the image.
@@ 4.1 Experimental Setup
Evaluation Metrics. We use bit per pixel (bpp) to measure the number of bits for latent representation compression. Two popular image quality assessment metrics, namely, PSNR and MS-SSIM [43], are used to evaluate the distortion between the reconstructed and original images. We also report the Bjontegaard Delta bitrate (BD-rate) [11] results to show average bitrate savings at equivalent distortion levels. In addition, both BD-PSNR and BD-MSSSIM are utilized to indicate average image quality improvements at constant bitrate.
Implementation Details. Leveraging CompressAI [8], we train our models with six different $\lambda$ values (256, 512, 1024, 2048, 4096, 8192 for the MSE metric; 8, 16, 32, 64, 128, 256 for the MS-SSIM metric). For MSE-optimized models, they are trained for 400 epochs with the Adam optimizer [26]. The batch size is set as $4$. The initial learning rate is $1e^{-4}$ and decayed by a factor of $2$ every $100$ epochs. To accelerate the experiments, both the image encoder and decoder utilize the pretrained single image compression model ELIC [23] provided by an open-source projecthttps://github.com/VincentChandelier/ELiC-ReImplemetation.git as the initial weights. For MS-SSIM evaluation, the MSE-optimized models are fine-tuned for $300$ epochs using the MS-SSIM distortion loss with the initial learning rate as $5e^{-5}$. During inference, we set the number of decoding steps as 8. Our experiments are conducted with NVIDIA V100 GPUs using PyTorch. For more details about the network structures, please refer to the supplementary material.
Benchmarks. We compare our CAMSIC with a variety of traditional and learning-based codecs. These competitive baselines are roughly split into four categories: single image compression (BPG [9], ELIC [23]), video compression (HEVC [38], VVC [12]), multi-view compression (MV-HEVC [39], LDMIC [50]) and stereo image compression (DispSIC [49], SASIC [45], BCSIC [27], ECSIC [44]). For BPG, we adopt the default x265 encoder without chroma subsampling to independently compress each image. By treating a stereo image pair as a two-frame video sequence, we run HM-18.0 [1] and VTM-23.0 [3] software with lowdelay P configuration and YUV444 format to evaluate the coding efficiency of HEVC and VVC. For MV-HEVC, HTM-16.3 [2] software is used to compress stereo images by setting a two-view intra mode. Unfortunately, it only supports YUV420 format, resulting in inferior compression performance. As for learning-based multi-view and stereo codecs except ECSIC, we implement them using the same training procedure. Since the current SOTA method ECSIC requires training on large-width visible stereo images to achieve the best results, we follow their open source libraryhttps://github.com/mwoedlinger/ecsic.git to train the ECSIC models. We refer readers to [44] for more training details about ECSIC.
@@
Methods BD-PSNR (dB) $\uparrow$ BD-rate (%) $\downarrow$ BD-MSSSIM (dB) $\uparrow$ BD-rate (%) $\downarrow$
HEVC [38] 0.499 -13.533 0.465 -14.453
VVC [12] 1.621 -37.213 1.538 -38.075
MV-HEVC [39] -1.693 61.915 -0.736 24.305
ELIC [23] 1.128 -25.196 1.198 -28.346
DispSIC [49] 0.402 11.865 2.948 -56.670
SASIC [45] -0.021 0.661 2.208 -43.415
BCSIC [27] 1.605 -35.891 3.355 -60.292
LDMIC [50] 1.600 -36.264 3.573 -64.398
ECSIC [44] 1.700 -36.665 3.274 -59.543
Proposed 2.019 -41.961 3.715 -64.519
@@
Methods BD-PSNR (dB) $\uparrow$ BD-rate (%) $\downarrow$ BD-MSSSIM (dB) $\uparrow$ BD-rate (%) $\downarrow$
HEVC [38] 0.103 -4.784 0.092 -4.260
VVC [12] 0.802 -29.125 0.669 -26.483
MV-HEVC [39] -1.274 77.154 -0.578 28.493
ELIC [23] 0.823 -29.748 0.807 -30.453
DispSIC [49] 0.617 -23.084 1.740 -46.916
SASIC [45] 0.231 -7.825 1.499 -39.693
BCSIC [27] 1.249 -39.933 2.228 -55.022
LDMIC [50] 1.282 -43.522 2.333 -59.167
ECSIC [44] 1.428 -43.551 2.149 -52.587
Proposed 1.440 -43.608 2.426 -56.773
Figure 5: Visualization comparison of our method against other codecs on the Cityscapes dataset. The best results are outlined in red color.
4.2 Experimental Results
@@
Compression performance. Fig. 4 shows the RD curves of these methods on Cityscapes and InStereo2K datasets. We also report BD-rate, BD-PSNR and BD-MSSSIM scores of each codec relative to BPG in Table 1. It is observed that the single image codec ELIC [23] achieves 25.196% and 29.748% bitrate saving in terms of PSNR on Cityscapes and InStereo2K datasets. By contrast, our CAMSIC has larger improvements, i.e., 41.961% and 43.608%. This demonstrates that introducing a stereo conditional entropy model to capture the correlations between stereo images effectively enables a more accurate estimation of the probability distribution.
@@
Without manually designed and complex linear or nonlinear transformations, our method outperforms most of these compression baselines, highlighting the importance of a powerful spatial-disparity entropy model. Specifically, a series of unidirectional codecs (HEVC [38], VVC [12], MV-HEVC [39], DispSIC [49], SASIC [45]) necessitate explicit warping to remove disparity redundancies. Notably, our CAMSIC surpasses VVC, the best unidirectional codec, across all bits per pixel (bpp) ranges on the InStereo2K and Cityscapes dataset. In addition, bidirectional codecs (BCSIC [27], ECSIC [44]) adopt bidirectional attention to reduce inter-view redundancies. Remarkably, even with a straightforward image encoder-decoder architecture, our CAMSIC still saves 8.512% and 0.629% bits over the SOTA method ECSIC in terms of PSNR on Cityscapes and InStereo2K datasets, respectively. Moreover, compared with the latest distributed codec LDMIC that adopts the joint decoding method based on global cross-attention mechanism, our proposed method improves by about 0.387dB and 0.195dB PSNR at the same bpp level on two datasets (Cityscapes and InStereo2K). These results affirm the superiority of our proposed entropy model in exploiting data correlations for enhancing compression performance.
Fig. 5 presents a subjective comparison of our proposed method against BCSIC, LDMIC, and ECSIC on two stereo image pairs from the Cityscapes dataset, where our method effectively restores the image details and exhibits superior reconstruction quality while consuming lower bits. Additional subjective comparisons for samples from both the Cityscapes and InStereo2K datasets are available in the supplementary material.
@@
Figure 6: Complexity of learning-based codecs on InStereo2K dataset. Both encoding and decoding time are evaluated by an NVIDIA V100 GPU.
@@
Coding Complexity. We compare the computation complexity of six learning-based multi-view and stereo codecs on InStereo2K dataset. The running time on GPU includes the execution latency of entropy coder. As illustrated in Fig. 6, our CAMSIC method is orders of magnitude faster than other auto-regressive approaches (DispSIC [49], BCSIC [27] and LDMIC [50]) by leveraging the parallel coding advantages of the MIM technique. Furthermore, by simplifying the stereo image encoding and decoding processes, our approach runs faster than SASIC [45] and ECSIC [44]. These results showcase a better trade-off between performance and speed among contemporary learning-based codecs.
4.3 Ablation Study
@@
Different Entropy Models. To demonstrate the superiority of our proposed entropy model, a set of ablation experiments are conducted. Fig. 7 presents the RD curves on the Cityscapes and InStereo2K datasets. Specifically, we replace our entropy model with the CNN-based ECSIC entropy model [44] and the content-irrelevant encoder-decoder Transformer entropy model presented in Section 3.2. It is observed that that the vanilla MIM-based Transformer entropy model outperforms the ECSIC entropy model, which is attributed to the Transformer’s inherent capability to capture long-range dependencies more effectively than CNN architectures. However, simply migrating the vanilla encoder-decoder Transformer architecture and the naive MIM technique to the entropy model results in inefficient prior information propagation. To address this challenge and fully unleash the potential of the Transformer, we introduce a novel content-aware MIM technique to design an advanced decoder-free Transformer entropy model. The improvements observed in Fig. 7 demonstrate that our content-aware MIM significantly surpasses the content-irrelevant MIM. This enhancement underscores the importance of bidirectional interactions between prior information and estimated tokens in improving compression performance. Through the strategic modifications, our method not only leverages the strengths of the Transformer architecture but also optimizes it for the specific demands of image compression, thereby achieving superior results.
@@
Operations for Token Interaction. In our framework, we enable relative position embedding and introduce identity embeddings to facilitate token interaction. To show the effects of these operations, we undertake a series of ablation experiments with BD-PSNR and BD-rate results. As shown in Table 2, compared with removing relative position embeddings (variant V2), adding relative position embeddings (variant V1) saves more bitrate, which indicates that providing the position cues for token interaction enables a more accurate estimation of probability distribution. Based on variant V1, we attach two identity embeddings to the input token sequence to form our final version, which further reduces the bitrate consumption. This underscores that explicitly distinguishing content-aware tokens and estimated tokens can improve coding gains.
@@
Table 2: Effect of different operations for token interaction. The first row represents our final solution and is set as the anchor to measure BD-PSNR and BD-rate. RPE means relative position embedding. ID indicates identitiy embedding.
Variants Cityscapes InStereo2K
BD-PSNR (dB) $\uparrow$ BD-rate (%) $\downarrow$ BD-PSNR (dB) $\uparrow$ BD-rate (%) $\downarrow$
Ours 0 0 0 0
(V1) w/o ID -0.104 2.739 -0.012 0.501
(V2) w/o RPE+w/o ID -0.140 3.700 -0.033 1.316
Figure 7: Ablation studies of different entropy models on Cityscapes and InStereo2K.
5 Conclusion

### v3

@@ Introduction
$\mathcal{L}$ $=\lambda D+R$ (7)
$=\lambda\sum_{v=1}^{2}d(\boldsymbol{x}_{v},\boldsymbol{\hat{x}}_{v})+\sum_{v=1}^{2}\big{(}R(\boldsymbol{\hat{y}}_{v})+R(\boldsymbol{\hat{z}}_{v})\big{)}$
where $d(\boldsymbol{x}_{v},\boldsymbol{\hat{x}}_{v})$ represents the distortion between the original image $\boldsymbol{x}_{v}$ and reconstructed image $\boldsymbol{\hat{x}}_{v}$ under a given metric such as MSE or MS-SSIM (Wang, Simoncelli, and Bovik 2003). $R(\boldsymbol{\hat{y}}_{v})$ denotes the estimated bitrate of the quantized representation $\boldsymbol{\hat{y}}_{v}$ and $R(\boldsymbol{\hat{z}}_{v})$ denotes the estimated bitrate of the hyper latent $\boldsymbol{\hat{z}}_{v}$. $\lambda$ is a hyper parameter that is used to balance the compression rate $R$ and distortion $D$.
Experiments
Experimental Setup
Dataset. We evaluate the coding efficiency of our proposed CAMSIC on two public stereo image datasets: (1) Cityscapes (Cordts et al. 2016): A dataset of urban outdoor scenes with distant views. It includes 5000 image pairs at a resolution of 2048$\times$1024 pixels. We divide these into 2975 pairs for training, 500 pairs for validation, and the remaining 1525 pairs for testing. (2) InStereo2K (Bao et al. 2020): A dataset of indoor scenes with close views. It contains 2060 image pairs at a resolution of 1080$\times$860 pixels. We allocate 2010 and 50 stereo image pairs for training and testing, respectively. We randomly crop the images into $256\times 256$ patches during training, and evaluate different codecs on full-resolution images. When the resolution of the source image is not supported, we apply the replicate padding to the right and bottom sides of the image.
@@
Evaluation Metrics. We use bit per pixel (bpp) to measure the number of bits for latent representation compression. Two popular image quality assessment metrics, namely, PSNR and MS-SSIM (Wang, Simoncelli, and Bovik 2003), are used to evaluate the distortion between the reconstructed and original images. We also report the Bjontegaard Delta bitrate (BD-rate) (Bjøntegaard 2001) results to show average bitrate savings at equivalent distortion levels. In addition, both BD-PSNR and BD-MSSSIM are utilized to indicate average image quality improvements at constant bitrate.
Implementation Details. Leveraging CompressAI (Bégaint et al. 2020), we train our models with 6 different $\lambda$ values (256, 512, 1024, 2048, 4096, 8192 for the MSE metric; 8, 16, 32, 64, 128, 256 for the MS-SSIM metric). For MSE-optimized models, they are trained for 400 epochs with the Adam optimizer (Kingma and Ba 2015). The batch size is set as $4$. The initial learning rate is $1e^{-4}$ and decayed by a factor of $2$ every $100$ epochs. To accelerate the experiments, both the image encoder and decoder utilize the pretrained single image compression model ELIC (He et al. 2022a). For MS-SSIM evaluation, the MSE-optimized models are fine-tuned for $300$ epochs using the MS-SSIM distortion loss with the initial learning rate as $5e^{-5}$. During inference, we set the number of decoding steps as 8. Our experiments are conducted with NVIDIA V100 GPUs using PyTorch.
Benchmarks. We compare our CAMSIC with a variety of traditional and learning-based codecs. These competitive baselines are roughly split into four categories: single image compression (BPG (Bellard 2014), ELIC (He et al. 2022a)), video compression (HEVC (Sullivan et al. 2012), VVC (Bross et al. 2021)), multi-view compression (MV-HEVC (Tech et al. 2015), LDMIC (Zhang, Shao, and Zhang 2022)) and stereo image compression (DispSIC (Zhai et al. 2022), SASIC (Wödlinger et al. 2022), BCSIC (Lei et al. 2022), ECSIC (Wödlinger et al. 2024)). For BPG, we adopt the default x265 encoder without chroma subsampling to independently compress each image. By treating a stereo image pair as a two-frame video sequence, we run HM-18.0 and VTM-23.0 software with lowdelay P configuration and YUV444 format to evaluate the coding efficiency of HEVC and VVC. For MV-HEVC, HTM-16.3 software is used to compress stereo images by setting a two-view intra mode. Unfortunately, it only supports YUV420 format, resulting in inferior compression performance. As for learning-based multi-view and stereo codecs except ECSIC, we implement them using the same training procedure. Since the current SOTA method ECSIC requires training on large-width stereo images to achieve the best results, we follow their open source library (Wödlinger et al. 2024) to train the ECSIC models.
@@
Methods BD-PSNR $\uparrow$ BD-rate $\downarrow$ BD-MSSSIM $\uparrow$ BD-rate $\downarrow$
HEVC 0.499dB -13.533% 0.465dB -14.453%
VVC 1.621dB -37.213% 1.538dB -38.075%
MV-HEVC -1.693dB 61.915% -0.736dB 24.305%
ELIC 1.128dB -25.196% 1.198dB -28.346%
DispSIC 0.402dB 11.865% 2.948dB -56.670%
SASIC -0.021dB 0.661% 2.208dB -43.415%
BCSIC 1.605dB -35.891% 3.355dB -60.292%
LDMIC 1.600dB -36.264% 3.573dB -64.398%
ECSIC 1.700dB -36.665% 3.274dB -59.543%
Proposed 2.019dB -41.961% 3.715dB -64.519%
@@
Methods BD-PSNR $\uparrow$ BD-rate $\downarrow$ BD-MSSSIM $\uparrow$ BD-rate $\downarrow$
HEVC 0.103dB -4.784% 0.092dB -4.260%
VVC 0.802dB -29.125% 0.669dB -26.483%
MV-HEVC -1.274dB 77.154% -0.578dB 28.493%
ELIC 0.823dB -29.748% 0.807dB -30.453%
DispSIC 0.617dB -23.084% 1.740dB -46.916%
SASIC 0.231dB -7.825% 1.499dB -39.693%
BCSIC 1.249dB -39.933% 2.228dB -55.022%
LDMIC 1.282dB -43.522% 2.333dB -59.167%
ECSIC 1.428dB -43.551% 2.149dB -52.587%
Proposed 1.440dB -43.608% 2.426dB -56.773%
Experimental Results
@@
Compression performance. Fig. 4 shows the RD curves of these methods on Cityscapes and InStereo2K datasets. We report BD-rate, BD-PSNR and BD-MSSSIM scores of each codec relative to BPG in Table 1. It is observed that the single image codec ELIC achieves 25.196% and 29.748% bitrate saving in terms of PSNR on Cityscapes and InStereo2K datasets. By contrast, our CAMSIC has larger improvements, i.e., 41.961% and 43.608%. This demonstrates that introducing a stereo conditional entropy model to capture the correlations between stereo images effectively enables a more accurate estimation of the probability distribution.
@@
Without manually designed and complex linear or nonlinear transformations, our method outperforms most of these compression baselines, highlighting the importance of a powerful spatial-disparity entropy model. Specifically, a series of unidirectional codecs (HEVC, VVC, MV-HEVC, DispSIC, SASIC) necessitate explicit warping to remove disparity redundancies. Notably, our CAMSIC surpasses VVC, the best unidirectional codec, across all bpp ranges on the InStereo2K and Cityscapes dataset. In addition, bidirectional codecs (BCSIC, ECSIC) adopt bidirectional attention to reduce inter-view redundancies. Remarkably, even with a straightforward image encoder-decoder architecture, our CAMSIC still saves more bits over the SOTA method ECSIC in terms of PSNR and MS-SSIM. Moreover, compared with the latest distributed codec LDMIC that adopts the joint decoding method based on global cross-attention mechanism, our proposed method improves by about 0.419dB and 0.158dB PSNR at the same bpp level on two datasets (Cityscapes and InStereo2K). These results affirm the superiority of our proposed entropy model in exploiting data correlations for enhancing compression performance.
@@
Figure 5: Complexity of learning-based codecs on the InStereo2K dataset. Both encoding and decoding time are evaluated by an NVIDIA V100 GPU.
@@
Coding Complexity. We compare the computation complexity of six learning-based multi-view and stereo codecs on InStereo2K dataset. The running time on GPU includes the execution latency of entropy coder. As illustrated in Fig. 5, our CAMSIC method is orders of magnitude faster than other auto-regressive approaches (DispSIC, BCSIC and LDMIC) by leveraging the parallel coding advantages of the MIM technique. Furthermore, by simplifying the stereo image encoding and decoding processes, our approach runs faster than SASIC and ECSIC. These results showcase a better trade-off between performance and speed among contemporary learning-based codecs.
Ablation Study
@@
Different Entropy Models. To demonstrate the superiority of our proposed entropy model, a set of ablation experiments are conducted. Table 2 presents the BDBR results on the Cityscapes and InStereo2K datasets. Specifically, we replace our entropy model with the CNN-based ECSIC, SASIC and BCSIC entropy models as well as the content-irrelevant encoder-decoder Transformer entropy model. Note that the vanilla MIM-based Transformer entropy model outperforms the ECSIC entropy model, which is attributed to the Transformer’s inherent capability to capture long-range dependencies more effectively than CNN architectures. However, simply migrating the vanilla encoder-decoder Transformer architecture and the naive MIM technique to the entropy model leads to inefficient prior information propagation, which makes the vanilla MIM entropy model inferior to the BCSIC entropy model with bidirectional view interaction. To address this challenge and fully unleash the potential of the Transformer, we introduce a novel content-aware MIM technique to design an advanced decoder-free Transformer entropy model. The improvements demonstrate that our content-aware MIM significantly surpasses the content-irrelevant MIM, which underscores the importance of bidirectional interactions between prior information and estimated tokens in improving compression performance.
@@
Operations for Token Interaction. In our framework, we enable relative position embedding and introduce identity embeddings to facilitate token interaction. To show the effects of these operations, we undertake a series of ablation experiments. As shown in Table 2, compared with removing relative position embeddings (variant V5), adding relative position embeddings (variant V6) saves more bitrate, which indicates that providing the position cues for token interaction enables a more accurate estimation of probability distribution. Based on variant V6, we attach two identity embeddings to the input token sequence to form our final version, which further reduces the bitrate consumption. It implies that explicitly distinguishing content-aware tokens and estimated tokens boosts token interaction.
@@
Table 2: Ablation studies for different components. The first row represents our final solution and is set as the anchor to measure BD-rate. RPE means relative position embedding. ID indicates identitiy embedding.
Variant Cityscapes InStereo2K
Ours 0% 0%
(V1) w/ ECSIC entropy model 48.89% 22.38%
(V2) w/ SASIC entropy model 36.16% 19.44%
(V3) w/ BCSIC entropy model 5.36% 5.76%
(V4) w/ Vanilla MIM entropy model 36.65% 19.48%
(V5) w/o RPE + w/o ID 3.700 1.316
(V6) w/o ID 2.739 0.501
Conclusion
@@
Acknowledgments
This work was supported by the General Research Fund (Project No. 16209622) from the Hong Kong Research Grants Council.
Appendix A Network Structure
Figure 6: The network structures of (a) hyperprior encoder, (b) hyperprior decoder, (c) prior fusion module with the details of the residual block shown in (d). “Conv (N, K, S)” represent the convolution operations with the number of output channels as $N$, the kernel size as $K\times K$ and the stride as $S$.
Figure 7: The network structure of (a) a Swin Transformer block and (b) entropy parameter module. W-MSA and SW-MSA are multi-head self attention modules with regular and shifted windowing configurations, respectively. LN means layer norm. MLP shorts for multi-layer perception. Linear (input size, output size).
Image Encoder and Decoder. The network structures of our encoder $E$ and decoder $D$ are the same as ELIC (He et al. 2022a), which stacks multiple residual bottleneck blocks and simplified attention modules. The number of channels in convolutional layers is set as 320.
Figure 8: The scheduled parallel decoding of MIM method. During inference, we decode $\hat{\boldsymbol{y}}_{v}$ in a few iterations. Each iteration recovers a portion of the most certain tokens with the smallest entropy based on the decoded ones.
Figure 9: Visualization comparison of our method against other codecs on the Cityscapes dataset.
Figure 10: Visualization comparison of our method against other codecs on the InStereo2K dataset.
Prior Generation Network. Fig. 6 shows the details of the hyperprior encoder/decoder and prior fusion module. After generating the hyper prior and disparity prior information, we apply a prior fusion module to produce the content-aware tokens $\boldsymbol{u}_{v}$.
Decoder-free Transformer. For the projection module, we use a linear layer to expand the dimensions from 320 to 768. As shown in Fig. 7 (a), we stack four post-norm Swin Transformer blocks (Liu et al. 2022b) to form our Swin Transformer encoder. The window size of W-MSA and SW-MSA is 8. The number of attention heads is 12 for every attention layer. We use two layers of MLP with the input size, hidden size, and output size as 768, 3072 and 768, respectively. As for the entropy parameter module, we apply three linear layers with Leakey ReLU activation function shown in Fig. 7 (b).
Appendix B Iterative Decoding Process
We follow the iterative decoding process in Chang et al. (2022) to generate an image during inference. Taking the content-aware masked image modeling (MIM) method as an example, we begin with all the tokens generated by prior information, i.e. $\boldsymbol{c}_{v}$. For ease of notation, let $\boldsymbol{u}_{v}^{0}=\boldsymbol{c}_{v}$. As shown in Fig. 8, the decoding process for iteration $k$ runs as follows:
Predict: Given the input tokens $\boldsymbol{u}_{v}^{k}$ of the decoder-free Transformer at the current iteration, our entropy model predicts the entropy of a token:
$H(\hat{\boldsymbol{y}}_{v}^{i,k})$ $=-\log_{2}\prod_{j=1}^{d}p(\mu_{v}^{i,j,k}),$ (8)
$p(\mu_{v}^{i,j,k})$ $=c_{v}^{i,j,k}(\mu_{v}^{i,j,k}+\frac{1}{2})-c_{v}^{i,j,k}(\mu_{v}^{i,j,k}-\frac{1}{2})$
where $d$ is the channel dimension of a token, $i$ is the position of the token, and $c_{v}^{i,j,k}(\cdot)$ is the cumulative function of Gaussian distribution $\mathcal{N}(\mu_{v}^{i,j,k},{\sigma_{v}^{i,j,k}}^{2})$.
Schedule: We compute the number of tokens $\left\lfloor\gamma(k)N-\gamma(k-1)N\right\rceil$ to be recovered at each step according to the scheduling function $\gamma(k)=\sin(k\cdot\frac{\pi}{2})$, where $N$ is the input token length.
Decode: We decode the selected tokens from bitstream with $(\boldsymbol{\mu}_{v}^{k},\boldsymbol{\sigma}_{v}^{k})$ and generate the input tokens $\boldsymbol{u}_{v}^{k+1}$ for iteration $k+1$.
Appendix C Visualizations
We conduct the visual comparison of our proposed method against BCSIC, LDMIC, and ECSIC. Several examples from Cityscapes and InStereo2K datasets are shown in Fig. 9 and 10. From these examples, we see that our approach achieves better rate-distortion performance.